# Unveiling Delay Effects in Traffic Forecasting: A Perspective from Spatial-Temporal Delay Differential Equations

### 02 Feb 2024 v1

@@ Abstract.
journalyear: 2024conference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeprice: 15.00isbn: 979-8-4007-0171-9/24/05
@@ 1. Introduction
Traffic forecasting is a fundamental research problem of Intelligent Transportation Systems (ITS) (Nagy and Simon, 2018; Yu et al., 2018; Ran and Boyce, 2012), which affects a variety of smart city applications (Choi et al., 2021; Song et al., 2020), such as trip planning (Choi et al., 2021; Fang et al., 2020; Aryandoust et al., 2019) accident prediction (Islam et al., 2022; Ju et al., 2023), and urban management (Chavhan and Venkataram, 2020; Yuan et al., 2021). Traffic flow forecasting aims to predict the future traffic flow based on historical data and underlying traffic networks.
Traffic flow forecasting is a challenging task due to the inherent spatial-temporal dependencies. Benefiting from the flourishing of deep learning, a large number of deep models have been proposed for traffic forecasting. In the temporal dimension, RNN-based models and their variants (Schuster and Paliwal, 1997) occupy the mainstream status, and temporal convolution networks (Lu et al., 2020) have also attracted much attention due to their superior computation efficiency. In the spatial dimension, considering that most traffic networks are non-Euclidean other than grid-partitioned, GNN-based methods (Yu et al., 2018; Li et al., 2018) beat CNN-based ones (Zhang et al., 2019) and become predominant owing to their strong ability to deal with graph-structured data. Extensive works combine the spatial module and the temporal module to achieve significant improvements, among which STGCN (Yu et al., 2018) and DCRNN (Li et al., 2018) and DSTAGN (Lan et al., 2022) are the representative.
@@
(2) The inherent continuity in traffic system is not well-explored. Existing methods mainly utilize RNNs (Schuster and Paliwal, 1997) or TCNs (Lu et al., 2020), which accept discrete observations as input, to capture the temporal dependencies. These methods are limited in terms of flexibility and applicability. Specifically, for the same traffic system, the required prediction horizon and resolution may vary across different applied scenarios, and the model needs to be retrained for each specific demand. Also, traffic data is notable for its inherent sparsity (Liu et al., 2020; Zhou et al., 2020). This sparsity arises due to the limited availability of traffic sensors, particularly in extensive road networks. For example, the sampling precision of sensors deployed within the traffic network may be at a 10-min interval. However, during the prediction phase, we aspire to attain finer-grained forecasting precision to enable rapid responses to events, such as travel time planning (Aryandoust et al., 2019) or traffic emergency management (Zhou et al., 2020).
@@
  • We propose a continuous output module, allowing us to accurately predict traffic flow at various frequencies, which provides more flexibility and adaptability to different scenarios.
  • We conduct experiments on six popular used datasets, in which results show that our model outperforms the SOTAs and exhibits competitive computational efficiency. Moreover, both quantitative and qualitative experiments are conducted to validate the concept of delay-aware module. Also, the flexibility validation shows the effectiveness of the continuous output module.
@@ 2.1. Traffic Flow Forecasting
A large body of research has been conducted on traffic flow forecasting in recent years. Traffic flow forecasting can be viewed as a spatial-temporal forecasting task leveraging spatial-temporal data collected by various sensors to predict future traffic conditions. In recent years, deep learning methods have dominated traffic flow forecasting issues, due to their superior ability to model complex spatial-temporal correlations. The models combining graph neural networks (GNN) (Kipf and Welling, 2016) and recurrent neural networks (RNN) (Schuster and Paliwal, 1997) are the representative. Specifically, DCRNN (Li et al., 2017) views the traffic flow as a diffusion process on a directed graph and utilizes GRU to capture the temporal features. STGCN (Yu et al., 2018) utilizes graph convolution and 1D convolution to capture spatial dependencies and temporal correlations respectively. ASTGCN(Guo et al., 2019) introduces an attention mechanism to capture the dynamics of spatial dependencies and temporal correlations. And AGCRN (Bai et al., 2020) captures node-specific spatial and temporal correlations automatically without a pre-defined graph. (Jiang et al., 2023) proposes a PDFormer model. While PDFormer mentions the concept of delay, its core mechanism involves utilizing attention to the historical time series, rather than explicitly utilizing delay in the propagation process. It implies that it still cannot capture the intricate delay information among graph vertices.
In general, despite their achieved success, all existing works are limited to the spatial-temporal stacking structure and ignore the delay effect, which deviates from the real situation of traffic.
@@ 2.2. Neural Differential Equations
Few pioneering works have been conducted in traffic forecasting with a neural differential equation framework. STGODE (Fang et al., 2021) first utilizes the neural ODE to transform graph convolution into a continuous version to acquire a larger spatial-temporal receptive field. STG-NCDE (Choi et al., 2021) adopts the neural CDE to deal with irregular-sampled time series. Despite the success have achieved, none of these works consider the delay effect. In this paper, we first extend the NDDE to multi-variable conditions for spatial-temporal modeling and cooperate with NCDE to construct continuous traffic signal evolution.
@@ 4.4. Continuous Output Module
Finally, as our objective function in the context of traffic flow forecasting, we employ the widely-used Huber loss (Huber, 1992), which is known for its robustness in handling outliers compared to the squared error loss.
@@
The whole flow of STDDE is presented in Algo 1 in the Appendix.
@@ 4.5.1. Connection to Existing Works
Then we analyze the generalizability of STDDE from two perspectives: 1) In the temporal dimension, where GRU and its variants (Fu et al., 2016) can be considered as special cases of STDDE when integration $\int\frac{dh_{i}(t)}{dt}$ is discrete. 2) In the spatial dimension, when all the time-delays $\tau_{ij}$ are set to zero, STDDE will degenerate to general GNNs (Hamilton et al., 2017; Ju et al., 2023).
@@ 5.2. Baselines
  • Spatio-Temporal Graph Convolution Models including STGCN (Yu et al., 2018), STSGCN (Song et al., 2020), DCRNN (Li et al., 2018), AGCRN (Bai et al., 2020), ASTGCN (Guo et al., 2019), FOGS (Rao et al., 2022). STGCN utilizes graph convolution and 1D convolution to capture spatial-temporal correlations. STSGCN utilizes multiple localized spatial-temporal subgraph modules to capture the spatial-temporal correlations directly. DCRNN integrates graph convolution into an encoder-decoder gated recurrent unit. AGCRN captures node-specific spatial and temporal correlations automatically without a pre-defined graph. ASTGCN utilizes spatial and temporal attention mechanisms to model their dynamics. DSTAGN constructs a dynamic graph instead of relying on a pre-defined static one. FOGS utilizes first-order gradients rather than specific flows, which effectively circumvent issues associated with fitting irregularly-shaped distributions.
@@ 5.3. Experimental Settings
Table 2. Performance comparison of baselines and proposed STGDDE on six popular used real-world traffic datasets.
@@ 5.6.1. Ablation Studies
To verify the effectiveness of different modules of STGDDE, we conduct the following ablation experiments on PeMS04 dataset and compare results with its corresponding variants.
  • v1 (STDDE-no-delay): In this variant, we take no account of the delay effect, and thus the model degenerates to an ODE model, to verify whether capturing time delay signal is contributing.
  • v2 (STDDE-zero-history): We set the history function of STDDE as zero function, to verify the necessity of learnable history states.
@@ A.1. Time Complexity Analysis
The time complexity of the model is primarily associated with the ODE module, and the complexity of ODE depends on the solving algorithm. In this study, we employ the Euler solver, so the complexity of aggregation is inversely proportional to the step length $\eta$. As we have T time slots in total, the aggregation needs to be calculated $\frac{T}{\eta}$ times. As the complexity of neighbor aggregation is $O(E)$, the algorithm’s complexity ultimately becomes $O(\frac{TE}{\eta})$.
A.2. Algorithm Pseudocode
The whole flow of STDDE is presented as follows.
Algorithm 1 STDDE
0:     Input series $\{X_{t-T+1:t,1},X_{t-T+1:t,2},\cdots,X_{t-T+1:t,N}\}$ for all $N$ nodes with series length $T$, graph $G$
0:     Output series $\{X_{t+1:t+T^{\prime},1},X_{t+1:t+T^{\prime},2},\cdots,X_{t+1:t+T^{\prime},N}\}$ for all nodes with series length $T^{\prime}$
1:  Calculate delay $\tau$ for all node pairs through MCC (12) or a learnable estimator;
2:  Generate continuous paths $X$ for all nodes through natural cubic spline;
3:  Generate the history function $\phi(t)$ based on input data series through Eqn.(10);
4:  Generate the hidden states and update the state flow through STDDE (15);
5:  Acquire outputs through another STDDE evolution (16);
A.3. Theoretical Analysis
@@
From theorem1, we see that $\mu_{p}(A)$ is easy to calculate for $p=1,\infty$ or estimate for $p=2$, which brings convenience for the analysis.
@@
where $h(t)=(h_{0}(t),h_{1}(t),\cdots,h_{N}(t)$ is a vector, $A_{0},A_{k}\in\mathbb{R}^{n\times n}$ are constant matrix, and $h(t_{\tau_{k}})=(h_{1}(t-\tau_{k1}),h_{2}(t-\tau_{k2}),\cdots,h_{N}(t-\tau_{kN})$.
@@
From eq (9), without considering the control signal, we have
@@
The lower bound is easy, and we can reformulate the upper bound as
@@
where $A_{k}$ is a matrix of which the $i$-th row contains at most one non-zero element of $c\alpha_{ij}$, and $H(t_{\tau_{k}})=(h_{1}(t_{\tau_{k1}}),h_{2}(t_{\tau_{k2}}),\cdots,h_{n}(t_{\tau_{kn}}))$ is the re-group of delayed signals. Moreover, $K$ is equal to the maximum of node degrees, because each edge appears only once in these matrices. ∎
@@
The second equation holds because there are at most one non-zero element in each row of $A_{k}$, and $c\alpha_{ij(k)}$ denotes the element in row $i$. Due to that $\alpha_{ij}$ is the normalized weight for graph convolution, we have
@@
the condition (33) holds. So, the lower bound and upper bound of the proposed DDE are both asymptotically stable, and they reach the same stable point (Yorke, 1970). As a result, the proposed DDE is asymptotically stable.

### 26 Feb 2024 v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645688isbn: 979-8-4007-0171-9/24/05ccs: Information systems Spatial-temporal systemsccs: Mathematics of computing Continuous functionsccs: Applied computing Transportation
@@ 1. Introduction
Traffic forecasting is a fundamental research problem of Intelligent Transportation Systems (ITS) (Nagy and Simon, 2018; Yu et al., 2018; Ran and Boyce, 2012; Fang et al., 2022), which affects a variety of smart city applications (Song et al., 2020a; Ju et al., 2024), such as trip planning (Choi et al., 2021; Fang et al., 2020) accident prediction (Islam et al., 2022; Ju et al., 2023), and urban management (Chavhan and Venkataram, 2020; Yuan et al., 2021). Traffic flow forecasting aims to predict the future traffic flow based on historical data and underlying traffic networks.
Traffic flow forecasting is a challenging task due to the inherent spatial-temporal dependencies. Benefiting from the flourishing of deep learning, a large number of deep models have been proposed for traffic forecasting. In the temporal dimension, RNN-based models and their variants (Schuster and Paliwal, 1997) occupy the mainstream status, and temporal convolution networks (Lu et al., 2020) have also attracted much attention due to their superior computation efficiency. In the spatial dimension, considering that most traffic networks are non-Euclidean other than grid-partitioned, GNN-based methods (Yu et al., 2018; Li et al., 2018; Long et al., 2020) beat CNN-based ones (Zhang et al., 2019) and become predominant owing to their strong ability to deal with graph-structured data. Extensive works combine the spatial module and the temporal module to achieve significant improvements, among which STGCN (Yu et al., 2018) and DCRNN (Li et al., 2018) and DSTAGN (Lan et al., 2022) are the representative.
@@
(2) The inherent continuity in traffic system is not well-explored. Existing methods mainly utilize RNNs (Schuster and Paliwal, 1997) or TCNs (Lu et al., 2020), which accept discrete observations as input, to capture the temporal dependencies. These methods are limited in terms of flexibility and applicability. Specifically, for the same traffic system, the required prediction horizon and resolution may vary across different applied scenarios, and the model needs to be retrained for each specific demand. Also, traffic data is notable for its inherent sparsity (Liu et al., 2020; Zhou et al., 2020). This sparsity arises due to the limited availability of traffic sensors, particularly in extensive road networks. For example, the sampling precision of sensors deployed within the traffic network may be at a 10-minute interval. However, during the prediction phase, we aspire to attain finer-grained forecasting precision to enable rapid responses to events, such as travel time planning (Aryandoust et al., 2019) or traffic emergency management (Zhou et al., 2020).
@@
  • We propose a continuous output module, allowing us to accurately predict traffic flow at various frequencies, which provides more flexibility to different scenarios.
  • We conduct experiments on six popular used datasets, in which results show that our model outperforms the SOTAs and exhibits competitive computational efficiency.
@@ 2.1. Traffic Flow Forecasting
A large body of research has been conducted on traffic flow forecasting in recent years. Traffic flow forecasting can be viewed as a spatial-temporal forecasting task leveraging spatial-temporal data collected by various sensors to predict future traffic conditions. In recent years, deep learning methods have dominated traffic flow forecasting issues, due to their superior ability to model complex spatial-temporal correlations. The models combining graph neural networks (GNN) (Kipf and Welling, 2016; Long et al., 2021a, b) and recurrent neural networks (RNN) (Schuster and Paliwal, 1997) are the representative. Specifically, DCRNN (Li et al., 2017) views the traffic flow as a diffusion process on a directed graph and utilizes GRU to capture the temporal features. STGCN (Yu et al., 2018) utilizes graph convolution and 1D convolution to capture spatial dependencies and temporal correlations respectively. (Jiang et al., 2023) proposes a PDFormer model. While PDFormer mentions the concept of delay, its core mechanism involves utilizing attention to the historical time series, rather than explicitly utilizing delay in the propagation process. It implies that it still cannot capture the intricate delay information among graph vertices. In general, despite their achieved success, all existing works are limited to the spatial-temporal stacking structure and ignore the delay effect, which deviates from the real situation of traffic.
@@ 2.2. Neural Differential Equations
Few pioneering works have been conducted in traffic forecasting with a neural differential equation framework. STGODE (Fang et al., 2021) first utilizes the neural ODE to transform graph convolution into a continuous version to acquire a larger spatial-temporal receptive field. STG-NCDE (Choi et al., 2021) adopts the neural CDE to deal with irregular-sampled time series. Despite the success have achieved, none of these works take the delay effect into consideration. In this paper, we first extend the NDDE to multi-variable conditions for spatial-temporal modeling and cooperate with NCDE to construct continuous traffic signal evolution.
@@ 4.4. Continuous Output Module
Finally, as our objective function in the context of traffic flow forecasting, we employ the widely-used Huber loss, which is known for its robustness in handling outliers compared to the squared error loss.
@@
The time complexity analysis of STDDE is presented in the Appendix.
@@ 4.5.1. Connection to Existing Works
Then we analyze the generalizability of STDDE from two perspectives: 1) In the temporal dimension, where GRU and its variants (Fu et al., 2016) can be considered as special cases of STDDE when integration $\int\frac{dh_{i}(t)}{dt}$ is discrete. 2) In the spatial dimension, when all the time-delays $\tau_{ij}$ are set to zero, STDDE will degenerate to general GNNs (Ju et al., 2023; Song et al., 2020b).
@@ 5.2. Baselines
  • Spatio-Temporal Graph Convolution Models including STGCN (Yu et al., 2018), STSGCN (Song et al., 2020a), DCRNN (Li et al., 2018), AGCRN (Bai et al., 2020), ASTGCN (Guo et al., 2019), FOGS (Rao et al., 2022). STGCN utilizes graph convolution and 1D convolution to capture spatial-temporal correlations. STSGCN utilizes multiple localized spatial-temporal subgraph modules to capture the spatial-temporal correlations directly. DCRNN integrates graph convolution into an encoder-decoder gated recurrent unit. AGCRN captures node-specific spatial and temporal correlations automatically without a pre-defined graph. ASTGCN utilizes spatial and temporal attention mechanisms to model their dynamics. DSTAGN constructs a dynamic graph instead of relying on a pre-defined static one. FOGS utilizes first-order gradients rather than specific flows, which effectively circumvent issues associated with fitting irregularly-shaped distributions.
@@ 5.3. Experimental Settings
Table 2. Performance comparison of baselines and proposed STDDE on six popular used real-world traffic datasets.
@@ 5.6.1. Ablation Studies
To verify the effectiveness of different modules of STDDE, we conduct the following ablation experiments on PeMS04 dataset and compare results with its corresponding variants.
  • v1 (STDDE-no-delay): We ignore the delay effect, and thus the model degenerates to an ODE model, to verify whether capturing the time delay signal is contributing.
  • v2 (STDDE-zero-history): We set the history function as zero to verify the necessity of learnable history states.
@@ A.1. Time Complexity Analysis
The time complexity of STDDE is mainly associated with the ODE module, and the complexity of ODE depends on the solving algorithm. In this paper, we employ the Euler solver, so the complexity of aggregation is inversely proportional to the step length $\eta$. As we have $T$ time slots, the aggregation needs to be calculated $\frac{T}{\eta}$ times. As the complexity of neighbor aggregation is $O(E)$, the algorithm’s complexity finally becomes $O(\frac{TE}{\eta})$.
A.2. Theoretical Analysis
@@
From theorem1, we see that $\mu_{p}(A)$ is easy to calculate for $p=1,\infty$, or estimate for $p=2$, which brings convenience.
@@
where $h(t)=(h_{0}(t),\cdots,h_{N}(t)$ is a vector, $A_{0},A_{k}\in\mathbb{R}^{n\times n}$ are constant matrix, and $h(t_{\tau_{k}})=(h_{1}(t-\tau_{k1}),\cdots,h_{N}(t-\tau_{kN})$.
@@
Without considering the control signal in Eqn. (9), we have
@@
The lower bound is easy, and we reformulate the upper bound as
@@
where $A_{k}$ is a matrix in which the $i$-th row contains at most one non-zero element of $c\alpha_{ij}$, and $H(t_{\tau_{k}})=(h_{1}(t_{\tau_{k1}}),\cdots,h_{n}(t_{\tau_{kn}}))$ is the re-group of delayed signals. $K$ denotes the maximum of node degrees, as each edge appears only once in the matrices. ∎
@@
The second equation holds because there is at most one non-zero element in each row of $A_{k}$, and $c\alpha_{ij(k)}$ denotes the element in row $i$. As $\alpha_{ij}$ is the normalized weight for graph convolution, we have
@@
the condition (33) holds. Thus the lower and upper bound of DDE are both asymptotically stable (Yorke, 1970). ∎